# HOLESOM: Constraining the Properties of Slowly-Accreting Massive Black Holes with Self-Organizing Maps

### v3

@@ 4.2 The Golden Test: Sgr A
We test HOLESOM on the archetypal case of a MBH accreting in ADAF mode: Sgr A. Although this MBH represents a perfect test case for our approach, we highlight two significant differences between Sgr Aand the typical MBH that we will target with our code: First, the mass of Sgr A, and, to some extent, its Eddington ratio, are known from a variety of independent observations, starting from the dynamics of stars in its immediate vicinity (see, e.g., Ghez et al. 2008) to recent measurements by the Event Horizon Telescope (Event Horizon Telescope Collaboration et al., 2022). The mass is $M_{\bullet}=4.0^{+1.1}_{-0.6}\times 10^{6}\,{\rm M_{\odot}}$, with an Eddington ratio of $\sim 10^{-7}$. Our methodology, based on only a few photometric data points, aims at broadly characterizing the mass and accretion regime of the MBH, and cannot compete with such accurate measurements. However, it is instructive to check whether our estimate is compatible with the true value for Sgr A. Second, Sgr Ais located in the Galactic center, and characterized by heavy obscuration by interstellar dust (see, e.g., Fritz et al. 2016), which makes observations in the optical and UV practically impossible. Targeting extragalactic MBHs, or off-centered wandering MBHs in the Milky Way, will avoid this issue. However, for this test, we feed the code only with radio, sub-millimeters, and X-ray data points for Sgr A.
@@
Radio observations at 5, 50, and 950 GHz are taken from Markoff et al. (2001); Yuan et al. (2003); Dodds-Eden et al. (2009); at 230 GHz from Bower et al. (2019), and the X-ray luminosity of Sgr Ain its quiescent state is from Baganoff et al. (2003). The input features used are the same as those in Table 1, with the exception that we use only the SED slope between 5 and 950 GHz, and use a new slope between radio (950 GHz) and X-ray. This additional slope helps track the synchrotron peak, as no information is employed for wavelengths in between. We also include photometric errors following the approach described in (La Torre et al., 2024). We generate 50 new input datasets by drawing new random luminosities from a Gaussian distribution centered on the true luminosity with a width equal to the error. Each of these 50 datasets is then projected onto the trained SOM. As described in Section 4, the ranges of $M_{\bullet}$ and $\,{f_{\rm Edd}}$ are determined from the 2D likelihood surface of these datasets on the SOM. Figure 4 shows the probability distribution function (PDF) of the derived $M_{\bullet}$ and $\,{f_{\rm Edd}}$, compared to the true values for Sgr A, and the corresponding SED models, which are consistent with the observed photometric points and their errors.
4.3 Disentangling MBHs from Stellar-Mass Objects
@@
Figure 5: Comparison between 300 models for MBH SEDs (orange lines), randomly drawn from the training set to cover all the physical parameter space, and the photometry for neutron star Sco X-1 (blue squares). Luminosities are normalized to the X-ray value.
Figure 1 suggests that slowly-accreting MBHs do not generally follow the well-known relations between radio and X-ray luminosities typical of stellar-mass compact objects, where $L_{R}\propto\sqrt{L_{X}}$ Gallo et al. (2003, 2018); Gusinskaia et al. (2020). Furthermore, as shown for the neutron star SCO X-1333The photometry for Sco X-1 was retrieved from this link. in Figure 5, the two populations of MBHs and stellar-mass compact objects have remarkably different SEDs.
We test whether HOLESOM can filter out stellar-mass compact objects, as they would not be projected on the same map trained with SEDs compatible with the MBH population. In principle, this applies when a test stellar-mass BH has complete photometric data across the employed feature set. In practice, the challenge arises with missing data. While in the previous sections we “filled-in” missing MHB data using luminosity or SED slope distributions from the MBH training set, now such distributions for the stellar-mass BH population are not available. Using MBH distributions to fill in missing stellar-mass BH data would make their SEDs more similar to those of MBHs, potentially blurring the distinction between the two populations.
[…]
In conclusion, HOLESOM effectively identifies and classifies MBHs based on their SEDs, projecting them into pixels that reflect their SEDs and parameters. In contrast, objects with significantly different SEDs, like stellar-mass BHs, are typically placed outside the map or at the edges, indicating they are outliers while still yielding parameters consistent with low masses and accretion closer to the Eddington limit.
@@ 5 Analytical Relations
Unlike stellar-mass BHs, which follow reasonably tight relations between radio and X-ray luminosities (see, e.g., Gallo et al. 2003), slowly-accreting MBHs do not exhibit such a strong correlation, as visually illustrated in Fig.1. However, we observe that these luminosities depend on $\,{f_{\rm Edd}}$ and $M_{\bullet}$. These insights drive us to seek analytical expressions tailored for slowly-accreting MBHs in a wide range of black hole masses and Eddington ratios. In the following, we present only those relations that effectively balance complexity and accuracy.
@@
$\ell_{m}=\alpha_{1}\ell_{r}^{2}+\beta_{1}\ell_{r}+\gamma_{1}$ (2)
@@
$\ell_{f}=\alpha_{2}\ell_{X}^{2}+\beta_{2}\ell_{X}+\gamma_{2}$ (3)
@@
Bottom: Eddington ratio versus luminosity in X-ray in erg///s. The red solid line and dashed and dotted lines are the same as in the top panel.
@@
The best equation resulting from symbolic regression is:
@@
$+(\gamma_{3}\ell_{m}+\delta_{3})\ell_{m}+\epsilon$
with the constants $\alpha_{3}$, $\beta_{3}$, $\gamma_{3}$, $\gamma_{3}$, and $\epsilon$ reported in Table 2. As shown in Figure 7, the predicted radio luminosity closely matches the true values, except for a few points in the $10^{27}-10^{29}$ erg/s range. This is due to the abrupt decrease of $M_{\bullet}$ at X-ray luminosity around $10^{38}-10^{39}$ erg/s, as seen in Fig. 1. We quantify the goodness of our prediction by computing the root mean square error (RMSE) and the mean absolute error (MAE), which are reported in Fig. 7.
@@ 6 Conclusions
Detecting MBHs accreting at extremely low rates in radiatively inefficient modes remains a significant challenge in astrophysics. These objects, which are easily confused/contaminated with other classes of sources, require advanced techniques for accurate identification and characterization. In this paper, we introduced HOLESOM, a ML tool based on the SOM algorithm. Trained with $\sim 20,000$ analytically-derived SEDs for ADAF accreting MBHs, HOLESOM  (i) automatically identifies and separates MBHs accreting in ADAF mode from other classes of objects, and (ii) efficiently characterize their physical properties, even with scarce data. We also used polynomial and symbolic regression to derive analytical relationships between radio/X-ray luminosities and black hole mass and Eddington ratio, providing additional insights into the behavior of MBHs radiating in ADAF mode.
@@
A key result of this work is HOLESOM’s ability to estimate mass and Eddington ratio ranges of MBHs with high accuracy, even in the presence of missing photometric data, which are recovered following La Torre et al. (2024). As expected, the process becomes more challenging when the available data is very scarce, particularly for MBHs accreting at very low Eddington ratios. In such cases, the accuracy of the mass and Eddington ratio estimates diminishes, but they are still consistent with their true values within uncertainties in the majority of cases. This highlights the importance of adding multi-wavelength data to improve the characterization of these objects. However, we demonstrated the robustness of our method by applying the model to the real-world case of Sgr A, where we retrieve distributions of $M_{\bullet}$ and $\,{f_{\rm Edd}}$ that were in excellent agreement with known values.
@@
We also showed that a promising aspect of HOLESOM is its ability to disentangle the photometric signatures of MBHs from those of stellar-mass BHs (e.g., X-ray binaries) in the same dataset. By projecting objects onto the trained SOM, MBHs are placed in regions corresponding to their expected parameter values, while objects with significantly different SEDs, such as stellar-mass BHs, are typically identified as outliers, assigned to the edges of the map, or not projected onto the map at all.
@@
In conclusion, HOLESOM provides a versatile and scalable framework for identifying and characterizing MBHs based on their SEDs. The usefulness of this method will become even more evident with large, multiwavelegth, low-z datasets. Once MBH candidates are identified in these surveys, additional multi-wavelength data will progressively hone in on the object’s physical properties, optimizing the effectiveness of HOLESOM and minimizing the impact of multiple missing photometric bands. As these surveys continue to grow, the ability to automatically identify MBHs and estimate their properties will be invaluable for advancing our understanding of the role these objects play in the evolution of galaxies.
@@
Acknowledgments: V.L.T. acknowledges support from the Smithsonian Astrophysical Observatory Visiting Fellowship, and the Kathryn A. McCarthy Graduate Fellowship in Physics administered by Tufts University. F.P. acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. This work was also supported by the Black Hole Initiative at Harvard University, funded by grants from the John Templeton Foundation and the Gordon and Betty Moore Foundation. This work used the FASRC cluster computer at Harvard University.
Software: SomPY (Moosavi et al., 2014b), PySR (Cranmer, 2023), numpy.polyfit.
@@ Appendix A Ranges of $\,{f_{\rm Edd}}$ and $M_{\bullet}$ in case of missing data
Figure 10: Range of Log $\,{f_{\rm Edd}}$ for test MBHs from the five test pixels (indicated in bold text) in the case of recovered missing multiple observations, as listed on the y-axis. The black cross and square for the cases ’TRUE’ and ’all’ are the same as in Figures 8-9.

### v4

@@ 4.2 The Golden Test: Sgr A
We test HOLESOM on the archetypal case of a MBH accreting in ADAF mode: Sgr A. Although this MBH represents a perfect test case for our approach, we highlight two significant differences between Sgr Aand the typical MBH that we will target with our code. First, the mass of Sgr A, and, to some extent, its Eddington ratio, are known from a variety of independent observations, starting from the dynamics of stars in its immediate vicinity (see, e.g., Ghez et al. 2008) to recent measurements by the Event Horizon Telescope (Event Horizon Telescope Collaboration et al., 2022). The mass is $M_{\bullet}=4.0^{+1.1}_{-0.6}\times 10^{6}\,{\rm M_{\odot}}$, with an Eddington ratio of $\sim 10^{-7}$. Our methodology, based on only a few photometric data points, aims at broadly characterizing the mass and accretion regime of the MBH, and cannot compete with such accurate measurements. However, it is instructive to check whether our estimate is compatible with the true value for Sgr A. Second, Sgr Ais located in the Galactic center and characterized by heavy obscuration by interstellar dust (see, e.g., Fritz et al. 2016), making optical and UV observations practically impossible. Targeting extragalactic MBHs, or off-centered wandering MBHs (Di Matteo et al., 2023) in the Milky Way, will avoid this issue. Therefore, for this test, we fill-in optical and UV data before projecting Sgr Aon the trained map.
@@
Radio observations at 5, 50, and 950 GHz are taken from Markoff et al. (2001); Yuan et al. (2003); Dodds-Eden et al. (2009); at 230 GHz from Bower et al. (2019); IR observations are taken from Melia & Falcke (2001); Genzel et al. (2003); Schödel, R. et al. (2007), and the X-ray luminosity of Sgr Ain its quiescent state is from Baganoff et al. (2003). The input features used are the same as those in Table 1, except that we use only the SED slope between 5 and 950 GHz and use a new slope between radio (950 GHz) and X-ray. This additional slope helps track the synchrotron peak, as no information is employed for wavelengths in between. We also include photometric errors and upper limits following the approach described in (La Torre et al., 2024). We generate 50 new input datasets by drawing new random luminosities from a Gaussian distribution centered on the true luminosity with a width equal to the error. In the case of upper limits, we retain only values lower than the true luminosity. Each of these 50 datasets is then projected onto the trained SOM. As described in Section 4, the ranges of $M_{\bullet}$ and $\,{f_{\rm Edd}}$ are determined from the 2D likelihood surface of these datasets on the SOM. Figure 4 shows the probability distribution function (PDF) of the derived $M_{\bullet}$ and $\,{f_{\rm Edd}}$, compared to the true values for Sgr A, and the corresponding SED models, which are consistent with the observed photometric points and their errors.
4.3 Identifying MBHs
@@
Figure 5: Left: Sources in a catalog are identified as MBH candidates (or not) based on their BMU distance. BMU distance distributions are shown for MBHs with a complete feature set (orange), MBHs with recovered 3, 5, 6, and 7 recovered missing features (shades of gray), Sgr A(black), and Sco X-1 (blue). See Sec. 4.3 for details. Right: Comparison between 300 models for MBH SEDs (orange lines), randomly drawn from the training set to cover all the physical parameter space, and the photometry for neutron star Sco X-1 (blue squares). Luminosities are normalized to the X-ray value.
With the new generation of telescopes increasing the volume of observational data, it is crucial to develop automated methods for systematically identifying MBH candidates before characterizing their properties and for follow-up observations. Currently, no standard automated approach exists for this task.
HOLESOM is the first attempt to systematically identify slowly-accreting MBH candidates from a multi-wavelength catalog, leveraging its training on MBH SEDs. Since ADAF MBHs are the only population the algorithm has been trained on, when new sources are projected on the map, HOLESOM effectively acts as a “filter”, retaining sources that resemble MBH templates while rejecting those that do not. Specifically, HOLESOM can: i) recognize objects with photometry similar to those in the training set by assigning them to SOM cells that best represent their observed SEDs, and ii) reject any source whose observed photometry differs significantly from the MBH training set, classifying it as a non-MBH (or a MBH accreting at substantial levels).
When a test object is projected onto HOLESOM, its identification is determined by the BMU distance (Sec. 3.1), which quantifies the Euclidean distance between the object’s observed data and the closest SOM cell (the BMU). Each BMU represents an MBH template. We compute BMU distances for 19,000 MBH models with complete input features, 2,000 MBH models with recovered missing features, and two real cases–an ADAF MBH and a non-MBH source. Because we account for photometric errors and missing data, this results in a distribution of BMU distances rather than a single value. As illustrated in Figure 5, BMU distance quantifies similarity to the MBH models making up the trained grid, allowing us to set identification thresholds:
  • BMU distance $<$ 2.5: The test object is confidently identified as an MBH.
  • BMU distance 2.5-7: The identification is uncertain, likely due to insufficient photometric coverage. Additional observations at currently missing wavelengths will refine classification.
  • BMU distance $>$ 7: The object is confidently classified as non-MBH, indicating it belongs to a different astrophysical population.
A challenge arises when test objects lack complete photometric coverage. In the previous section, we fill in missing values using luminosity or SED slope distributions from the MBH training set. However, such distributions are not available for non-MBH objects. Applying MBH-based priors to non-MBH objects would make their SEDs appear more MBH-like, potentially blurring the distinction between the populations and reducing classification accuracy. Despite this challenge, we test HOLESOM’s identification capability with real-world examples and find that the method remains effective.
To evaluate HOLESOM’s performance, we project Sco X-1, a stellar-mass compact object, onto the map. As shown in Figure 5 (right), Sco X-1 has a remarkably different SED from those of MBHs accreting in ADAF mode. Of the 10 input features used for training (Table 1), six are missing. We only have the radio luminosity at 5 GHz, the X-ray luminosity, and the IR luminosities at 25.5, 8.6, and 1.5 $\mu$m, i.e., five bands out of the 10 needed to construct the full feature set. Despite recovering missing data using MBH priors (which could make Sco X-1’s SED more similar to MBHs), Sco X-1 still exhibits a high BMU distance $\gtrsim$ 7 (Figure 5 left). Note that MBHs with even more missing features still maintain lower BMU distances (0-5). This confirms that HOLESOM correctly rejects Sco X-1 as an MBH candidate, placing it at the edge of the map (i.e., in the SOM location [18,0]), where outliers are typically located (see, e.g., Davidzon et al. 2019). Conversely, when we project Sgr Aonto the map (Sec. 4.2), its BMU distance remains below 1.8 (Figure 5 left), reinforcing its identification as an ADAF MBH with high confidence, even with missing data.
By implementing the BMU distance threshold, HOLESOM provides a robust, automated framework for identifying ADAF MBH candidates in large datasets. Even when dealing with missing data, the method remains effective, ensuring reliable discrimination between MBH candidates and other astrophysical populations.
5 How to use HOLESOM with new observations
The population of slowly accreting MBHs remains elusive due to the challenge of detecting them. Only a few candidates have been observed, often with sparse multi-wavelength data (e.g., Nguyen et al., 2019; Reines, 2022) . Moreover, even knowing where to look for such objects is difficult. This work introduces HOLESOM as a tool to identify slowly accreting MBH candidates in large multi-wavelength catalogs containing diverse astrophysical sources. HOLESOM can also characterize their physical properties $M_{\bullet}$ and $\,{f_{\rm Edd}}$. Once identified, these candidates can benefit from follow-up observations to refine their parameters and spectroscopy to investigate their nature in detail. While this paper presents and validates the methodology using specific examples (Sgr Aand Sco X-1), a large-scale application to compile a catalog and project it onto HOLESOM is planned for future work.
Figure 6: Schematic describing the procedure that HOLESOM uses to identify MBH candidates and characterize their $M_{\bullet}$ and $\,{f_{\rm Edd}}$ from multi-wavelength catalogs.
In this section, we outline how the pre-trained SOM can be applied to photometric data of a new source, as schematically illustrated in Figure 6. The process involves projecting the available photometry onto the trained map using the same input features used during training. However, MBHs often have incomplete photometric coverage. In this case, we first recover missing observations at the required frequencies, as detailed in Section 4.1. Now, with a full set of input features, the source is projected onto the map, and the BMU distance is computed. As detailed in Section 4.3, based on this distance, the source is classified into one of three categories: i) high-confidence MBH candidate, ii) possible MBH candidate requiring additional observations for confirmation, and iii) not an MBH, likely belonging to a different astrophysical population. Once a candidate or possible candidate is identified, HOLESOM provides an estimate, with uncertainties, of $M_{\bullet}$ and $\,{f_{\rm Edd}}$, which could be refined with photometric or spectroscopic follow-up observations.
@@ 6 Analytical Relations
Unlike stellar-mass BHs, which follow reasonably tight relations between radio and X-ray luminosities (see, e.g., Gallo et al. 2003), slowly-accreting MBHs do not exhibit such a strong correlation, as visually illustrated in Fig.1. However, these luminosities depend on $\,{f_{\rm Edd}}$ and $M_{\bullet}$. These insights drive us to seek analytical expressions tailored for slowly-accreting MBHs in a wide range of black hole masses and Eddington ratios. In the following, we present only those relations that effectively balance complexity and accuracy.
@@
$\ell_{m}=\alpha_{1}\ell_{r}^{2}+\beta_{1}\ell_{r}+\gamma_{1}\,,$ (2)
@@
$\ell_{f}=\alpha_{2}\ell_{X}^{2}+\beta_{2}\ell_{X}+\gamma_{2}\,,$ (3)
@@
The best equation from symbolic regression is:
@@
$+(\gamma_{3}\ell_{m}+\delta_{3})\ell_{m}+\epsilon\,,$
with constants $\alpha_{3}$, $\beta_{3}$, $\gamma_{3}$, $\gamma_{3}$, and $\epsilon$ reported in Table 2.
As shown in Figure 8, the predicted radio luminosity closely matches the true values, except for a few points in the $10^{27}-10^{29}$ erg/s range. This is due to the abrupt decrease of $M_{\bullet}$ at X-ray luminosity around $10^{38}-10^{39}$ erg/s, as seen in Fig. 1. In Fig. 8, we quantify the goodness of our prediction by computing the root mean square error (RMSE) and the mean absolute error (MAE).
@@ 7 Conclusions
Detecting MBHs accreting at extremely low rates in radiatively inefficient modes remains a significant challenge in astrophysics. These objects, which are easily confused with other classes of sources, require advanced techniques for accurate identification and characterization.
@@
In this paper, we introduced HOLESOM, a machine learning tool based on the SOM algorithm. Trained with $\sim 20,000$ analytically-derived SEDs for ADAF accreting MBHs, HOLESOM  (i) automatically identifies and separates MBHs accreting in ADAF mode from other classes of objects, and (ii) efficiently characterize their physical properties, even with scarce data. We also used polynomial and symbolic regression to derive analytical relationships between radio/X-ray luminosities and black hole mass and Eddington ratio, providing additional insights into the behavior of MBHs radiating in ADAF mode.
@@
A key result of this work is HOLESOM’s ability to estimate mass and Eddington ratio ranges of MBHs with high accuracy, even in the presence of missing photometric data, which are recovered following La Torre et al. (2024). As expected, the process becomes more challenging when the available data is scarce, particularly for MBHs accreting at very low Eddington ratios. In such cases, the accuracy of the mass and Eddington ratio estimates diminishes, but they are still consistent with their true values within uncertainties in the majority of cases. This highlights the importance of adding multi-wavelength data to improve the characterization of these objects. However, we demonstrated the robustness of our method by applying the model to the real-world case of Sgr A, where we retrieved distributions of $M_{\bullet}$ and $\,{f_{\rm Edd}}$ that were in excellent agreement with known values.
@@
We also showed that a promising aspect of HOLESOM is its ability to recognize the photometric signatures of MBHs while rejecting different astrophysical sources, based on their BMU distance. By projecting objects onto the trained SOM, MBHs exhibit low BMU distances ($<2.5$) and are placed in regions corresponding to their expected parameter values. In contrast, objects with significantly different SEDs, such as stellar-mass BHs, exhibit considerable high (i.e., $>$7) BMU distances and are identified as outliers, assigned to the edges of the map, or not projected onto the map at all.
@@
In conclusion, HOLESOM provides a versatile and scalable framework for identifying and characterizing MBHs based on their SEDs. The capabilities of this method will become even more evident with large, multi-wavelength, low-z datasets. As part of our future work, we will apply HOLESOM to a multi-wavelength, low-z catalog to identify slowly-accreting MBH candidates in the local Universe and provide preliminary estimates of their mass and Eddington ratio, constructing a first attempt at a mass and Eddington ratio distribution functions. Once MBH candidates are identified, additional multi-wavelength data will progressively hone in on the object’s physical properties, optimizing the effectiveness of HOLESOM and minimizing the impact of multiple missing photometric bands. As these surveys continue to grow, the ability to automatically identify MBHs and estimate their properties will be invaluable for advancing our understanding of the role these objects play in the evolution of galaxies.
@@
Acknowledgments: The authors are thankful to the anonymous referees for their thoughtful and constructive feedback, which greatly improved this manuscript. V.L.T also thanks Dr. Kaitlin Gili for insightful discussions on supervised and unsupervised methods during the writing of this paper. V.L.T. acknowledges support from the Smithsonian Astrophysical Observatory Visiting Fellowship and the Kathryn A. McCarthy Graduate Fellowship in Physics administered by Tufts University. F.P. acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. This work was also supported by the Black Hole Initiative at Harvard University, funded by grants from the John Templeton Foundation and the Gordon and Betty Moore Foundation. This work used the FASRC cluster computer at Harvard University.
Software: SomPY (Moosavi et al., 2014), PySR (Cranmer, 2023), numpy.polyfit.
@@ Appendix A Ranges of $\,{f_{\rm Edd}}$ and $M_{\bullet}$ in case of missing data
Figure 11: Range of Log $\,{f_{\rm Edd}}$ for test MBHs from the five test pixels (indicated in bold text) in the case of recovered missing multiple observations, as listed on the y-axis. The black cross and square for the cases ’TRUE’ and ’all’ are the same as in Figures 9-10.